OpenStudy (anonymous):

Find the slope of the line through (3, –6) and (5, 8). (1 point)

OpenStudy (anonymous):

use the slope formula, that is to say m=(y2-y1/x2-x1)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Then if you want to write the equation of that particular line, use the formula: y-y1=m(x-x1)

OpenStudy (anonymous):

(-6-8)/(3-5) = -14/-2 = 7

OpenStudy (anonymous):

the formule is that : y2-y1 / x2-x1 then its 8-(-6) / 5-3 = 14/2 = 7